# Towards Scalable Exact Machine Unlearning Using Parameter-Efficient Fine-Tuning

### v1

@@ A.1 Proof of Lemma 1
This lemma states that BMS is optimal and selects diverse sequences. Two sequences are considered diverse if no element appears at the same position, e.g., $(S_{1},S_{2},S_{3},S_{4})$ and $(S_{2},S_{3},S_{4},S_{1})$. Since we want diverse sequences, BMS performs perfect matching, ensuring that no two elements appear in the same position. Therefore, in this proof, we show that perfect matching exists at all levels, $[1,L]$, in Algorithm 1.
@@
$|W|\leq N_{G}(W),$ (3)
where $N_{G}(\cdot)$ is the neighbourhood defined using the graph, $G$.
Since perfect matching selects a unique element in $V^{\prime}$, therefore no element is repeated at the same position in the BMS output. This happens for all iterations, BMS is optimal. ∎
A.2 Proof of Lemma 2
In S3T, when $B\geq L$ then all slices $\{1,\ldots,L\}$ must occupy the topmost position in one of the trained models (an example is shown in Figure 4). This implies that for S3T to encounter service failure, all slices in all shards must be affected by deletion requests. Therefore, considering deletion requests affect all slices uniformly we need to compute the expected time till all slices are affected. This setup is similar to the coupon collector problem [7].
Proof.
The deletion rate is the expected number of requests to delete all $mL$ slices ($m$ shards, $L$ slices per shard). Using linearity of expectation, the deletion rate or total time can be written as:
@@ A.2 Proof of Lemma 2
$=\mathbb{E}[t_{1}]+\ldots+\mathbb{E}[t_{mL}],$ (5)
@@
$p_{i}=\frac{\{mL-(i-1)\}s_{b}}{N-r}=\frac{N-(i-1)s_{b}}{N-r}=\frac{1-(i-1)/mL}{1-r/N}\geq\frac{mL-(i-1)}{mL},$ (6)
@@
$=mL\left(\frac{1}{1}+\frac{1}{2}+\ldots+\frac{1}{mL}\right)$ (8)
$=mL.H_{mL}$ (9)
$=mL\log(mL)+\gamma mL+\frac{1}{2}+O\left(\frac{1}{mL}\right),$ (10)
@@
This proves the second portion of the lemma: $\delta(\mathrm{SISA})\sim O(mL\log m)$. ∎
A.3 Proof of Lemma 3
@@
In S3T, each model is trained on $B$ such sequences. Therefore, the probability that all $B$ sequences are affected is:
@@
However, the above result suggests that the probability can be increased indefinitely by increasing $B$, which is not accurate because there are only $\binom{L}{k}$ possible sequences. To correct this, we show:
@@
where $B^{\prime}=\min\left\{B,\binom{L}{k}\right\}$. This proves the lemma. ∎
@@ Appendix B Implementation Details
B.1 Bipartite-matching based Sequence Selection
We describe the details of the Bipartite-matching based selection (BMS) algorithm in Algorithm 1. The objective of BMS is to select a set of $B$ diverse permutation sequences of length $L$ (slice count). This algorithm starts by selecting $L$ different starting elements for the sequences in Line 4. Then, BMS iteratively selects the next element within each sequence. This involves constructing a bipartite graph between the last elements of the sequences seen so far and the next set of elements. The edges incident on feasible next elements (elements not seen in a permutation sequence so far) as shown in Line 11. We perform bipartite graph matching on the graph, $G$. Based on the matching, we select the next element for each permutation sequence (Line 17). BMS continues this process till all sequences $o\in\mathcal{O}$ have $L$ elements. Based on the budget, we output $B$ sequences from $\mathcal{O}$.
For Bipartite graph matching, Hopcroft–Karp algorithm [30] provides an efficient algorithm with time complexity $O(|E|\sqrt{|V|})$, where $|E|,|V|$ are the number of edges and vertices respectively. Using this algorithm, the overall time complexity of BMS is $O(L^{2.5})$.
@@ B.1 Bipartite-matching based Sequence Selection
11:                  if $v^{\prime}\notin o$ then $G.\mathrm{add\_edge}(v,v^{\prime})$ // check for feasibility
12:              end for
@@
14:         $(V,V^{\prime})=\mathrm{PerfectMatching}(G)$ // compute bipartite perfect matching [30]
15:         for $o\in\mathcal{O}$ do
16:              // select vertex associated with each sequence from perfect matching
17:              ${v^{*}}=\{v^{\prime}|v^{\prime}\in V^{\prime}\land v=o.\mathrm{pop}()\}$
18:              $o=o\cup{v^{*}}$ // add to sequence
19:         end for
@@
21:     $\mathcal{O}_{B}=\{o_{1},\ldots,o_{B}\}\in\mathcal{O}$ // select any $B$ sequences from $\mathcal{O}$
22:     return $\mathcal{O}_{B}$
23:end function
B.2 S3T Training Procedure
In this section, we provide an outline for the training procedure within the S3T framework in Algorithm 2. S3T proceeds by dividing the entire dataset into $m$ shards. Each shard is further divided into $L$ disjoint slices. Based on the budget $B$, we obtain the permutation sequences to perform slice-wise training. We train each model $f$ on a unique sequence of slice sequence, $\pi(\mathcal{S})$. We return the complete set of models, $\mathcal{F}$. During inference, the user selects the best performing model within each shard and deploys an ensemble of those models to production.
@@ B.2 S3T Training Procedure
5:     if $B\leq L$ then
6:         $\Pi=\mathrm{BMS}(L,B)$ // perform BMS selection
7:     else
8:         $\Pi=\mathrm{ConditionalSampling}(L,B)$ // Conditional Sampling based selection
9:     end if
10:     for $\pi\in\Pi$ do
11:         $\mathcal{S}_{\pi}=\pi(\mathcal{S})$ // order the slices according to the permutation $\pi$
12:         $f=\mathrm{SliceWiseTraining}(\mathcal{S}_{\pi})$
13:         $\mathcal{F}=\mathcal{F}\cup f$
14:     end for
15:end for
16:return $\mathcal{F}$
@@
$\mathcal{F}=\mathcal{F}\setminus\overline{\mathcal{F}}$ // introduce updated models
@@ C.1 Experimental Setup
We perform all experiments using PyTorch [47] and Huggingface [64] framework. Our experiments were run on NVIDIA A6000 GPUs. In Table 2, we report the common set of hyperparameters for S3T fine-tuning experiments. All hyperparameters were set using a grid search with the Weights & Biases framework. We use an AdamW optimizer with the corresponding learning rates for each dataset (reported in Table 2). During fine-tuning of the models, we perform full-precision training for all settings except instruction tuning where we use 8-bit training.
@@
C.2 Additional Experiments
@@
Performance Degradation with Slice Deletion. In Figure 9, we report the relative performance drop with an increasing number of slices affected by deletion requests. For vision datasets in Figure 9 (left), we only observe a small drop in performance (<5%). For instruction tuning in Figure 9 (middle), we observe a negligible performance drop for most datasets except MMLU & ARC, which are more challenging open-ended QA benchmarks. For text classification in Figure 9 (right), we observe an increased drop in performance for a few of the datasets. We hypothesize that this occurs because the RoBERTaLARGE (used in text classification tasks) is a relatively weaker pre-trained model compared to Llama and ViT, which may explain this behavior. We also observe that the performance can vary based on the task and overall it is dependent on both the task and the model being fine-tuned. For unlearning, a lower performance drop is better as it suggests that we can continue using the same model without retraining for a longer time.
Figure 9: Relative performance drop with an increasing number of deleted slices. We observe a relatively small performance drop for image classification and instruction tuning tasks, while noticing a considerable drop for few of the text classification datasets.
@@
Layer Allocation per Slice. This experiment aims to answer the question: how many slices can we pack into a model while still achieving good performance? We conduct an experiment where we allocate different numbers of PEFT layers per slice and also vary the number of slices. We report the results on RTE dataset using RoBERTaLARGE in Figure 10. We observe that the performance is poor when the number of layers per slice is low. For example, when the number of layers per slice = 1, the performance improves as we increase the number of slices but again drops when the slice count is too high. This shows that it may not be feasible to train the model using a large number of slices without a drop in performance. Overall, the performance variation depends on the underlying task and model, so the developer should select these hyperparameters based on their performance requirements.
Sequence-based Training. In this experiment, we compare S3T’s performance with sequential training. Sequential training involves training the entire PEFT model on a sequence of slices. This was used in [35] to improve the retraining efficiency of SISA. In Figure 11, we report the performance of S3T and sequential training using ViTBASE on CIFAR-10 and CIFAR-100 datasets. We observe that the performance of S3T gradually increases as it is trained on more slices. Overall, we observe that S3T achieves quite similar or outperforms sequential training. It is important to note that S3T trains a significantly smaller number of parameters ($L$ times reduction compared to sequential training) but still achieves competitive performance.
@@
Appendix D Limitations
In this paper, we present a novel exact unlearning framework, S3T. S3T improves the deletion rate of existing unlearning systems at the cost of additional offline training. The training process can be time-consuming if we are fine-tuning larger models and have a higher performance requirement (thereby high budget $B$). However, this is an inherent tradeoff between offline training and losing revenue due to re-training costs. The developer should adjust their budget according to the tradeoff in their specific application.
Appendix E Broader Impact
We present a scalable approach to perform exact unlearning in production systems. We hope that this framework will be adopted by organizations and enable a cost-effective way to ensure the privacy requests of various users. In general, unlearning systems are susceptible to attacks where the adversary may design deletion requests in a way to modify the model behaviour according to their needs. Therefore, it is important to ensure that the deletion requests executed on the unlearning system are not malicious. However, this is a limitation of unlearning systems in general and not specific to our proposed framework, S3T. Future works can focus on the identification of malicious unlearning requests.

### v2

@@ A.1 Proof of Lemma 1
In S3T, since the sequences are selected to be diverse (Figure 4), the topmost slice in each sequence is different. Therefore, we have $B^{\prime}=\min(B,L)$ different slices at the topmost position. This implies that for S3T to encounter service failure, a total of $mB^{\prime}$ slices must be affected by deletion requests, where $m$ is the number of shards. Considering deletion requests affect all slices uniformly we need to compute the expected time till all slices are affected. This setup is similar to the coupon collector problem (Blom et al., 1994).
@@
The deletion rate is the expected number of requests to delete all $mB^{\prime}$ slices ($m$ shards, $B^{\prime}$ unique slices per shard). Using linearity of expectation, the deletion rate or total time is:
@@
$=\mathbb{E}[t_{1}]+\ldots+\mathbb{E}[t_{mB^{\prime}}],$ (5)
@@
$p_{i}$ =$=\frac{\{mB^{\prime}-(i-1)\}s_{b}}{N-r}$ (6)
$p_{i}$ $=\frac{N\min(B/L,1)-(i-1)s_{b}}{N-r}$ (7)
$p_{i}$ $=\frac{\min(B/L,1)-(i-1)/(mL)}{1-r/N}$ (8)
$\Rightarrow p_{i}$ $\geq\frac{mB^{\prime}-(i-1)}{mL},$ (9)
@@
$=mL\left(\frac{1}{1}+\frac{1}{2}+\ldots+\frac{1}{mB^{\prime}}\right)$ (11)
$=mL.H_{mB^{\prime}}$ (12)
$=mL\log(mB^{\prime})+\gamma mL+\frac{1}{2}+O\left(\frac{L}{mB^{\prime}}\right),$ (13)
@@
This proves the second part of the lemma: $\delta(\mathrm{SISA})\sim O(mL\log m)$. ∎
A.2 Proof of Lemma 2
@@
In S3T, each model is trained on $B$ such sequences. Therefore, we need to compute the probability that at least one sequence is better than $F(k)$, which is:
@@
However, the above result suggests that the probability can be increased indefinitely by increasing $B$. This is inaccurate because to maintain a performance of $F(k)$ there has to be at least one prefix of length $k$ that has not been affected. Since there only $P(L,k)=\frac{L!}{(L-k)!}$ permutations of length $k$ extending the budget beyond $P(L,k)$ doesn’t work. Therefore, the correct probability is:
@@
where $B^{\prime}=\min\left\{B,P(L,k)\right\}$.
Taking the difference between Eq. 20 and Eq. 18 and setting $\alpha=(1-k/L)^{r}$, we get:
[…]
$=(1-\alpha)\{1-(1-\alpha)^{B^{\prime}-1}\}$
$=\zeta(1-\zeta^{B^{\prime}-1}),$ (21)
where $\zeta=1-\alpha=1-(1-k/L)^{r}$. This completes the proof. ∎
Discussion. In the above proof, we assumed that $B$ sequences are randomly selected. In practice, we select diverse sequences using the iterative cyclic rotation algorithm. However, deriving a closed-form theoretical performance bound for the sequences generated using cyclic rotation is non-trivial. Intuitively, we expect the performance to be better as selecting diverse sequences means that the probability of a length-$k$ prefix getting affected is reduced. We leave the theoretical proof of these results to future work.
A.3 Proof of Lemma 3
This lemma states that BMS selects the most diverse sequences. First, we revisit our definition of sequence diversity. Two sequences are considered diverse if no element appears at the same position, e.g., $(S_{1},S_{2},S_{3},S_{4})$ and $(S_{2},S_{3},S_{4},S_{1})$. Since we want diverse sequences, BMS performs maximum weight perfect matching, ensuring that no two elements appear in the same position. Therefore, in this proof, we show that perfect matching exists at all levels, $[1,L]$, in Algorithm 2.
Proof.
$|W|\leq N_{G}(W),$ (22)
where $N_{G}(\cdot)$ is the neighbourhood defined using the graph, $G$.
Since perfect matching selects a unique element in $V^{\prime}$, therefore no element is repeated at the same position in the BMS output. Therefore, BMS generates diverse sequences. ∎
@@ Appendix B Implementation Details
Algorithm 1 Iterative Cyclic Rotation
1:function CyclicPermutation(Permutation $P$)
2:     ${}\mathcal{R}=\{\}$ // set of all rotations
3:     for $i\in\{1,\ldots,P.\mathrm{length}\}$ do
4:         $\mathcal{R}=\mathcal{R}\cup P$
5:         $P=\mathrm{rotateRight}(P)$
6:     end for
7:     return $\mathcal{R}$
8:end function
9:
10:function IterativeCyclicRotation(Slice count: $L$, Budget: $B$)
11:     $\mathcal{O}=\mathrm{CyclicPermutation}([1,\ldots,L])$ // initialize the set with $L$ cyclic permutations
12:     $n_{\mathrm{iter}}=0$ // set the iteration count
13:     while $|\mathcal{O}|<B$ do
14:         $n_{\mathrm{iter}}=n_{\mathrm{iter}}+1$
15:         for $o\in\mathcal{O}$ do // iteratively expand each of the existing permutation
16:              $\mathrm{prefix},\mathrm{suffix}=o[:n_{\mathrm{iter}}],o[n_{\mathrm{iter}}:]$ // set the prefix and suffix
17:              // rotate the suffix with same prefix
18:              $\mathcal{P}=\{\mathrm{prefix}\cup p\;\text{for }p\in\mathrm{CyclicPermutation}(\mathrm{suffix})\}$
19:              $\mathcal{O}=\mathcal{O}\cup\mathcal{P}$
20:         end for
21:     end while
22:     return $\mathcal{O}[:B]$ // return $B$ output permutations
23:end function
B.1 Iterative Cyclic Rotation
In the setting, where the deletion probabilities are uniform we use the cyclic rotation algorithm to select diverse permutations. In Figure 3 (middle), we observe that we can easily generate $L$ diverse sequences using cyclic permutation of the original sequence. We can iteratively expand on this idea to generate more sequences when the budget is $B>L$. An illustration of the different sequences selected using iterative cyclic rotation for different budgets is shown in Figure 10.
We present the pseudocode of the iterative cyclic rotation mechanism in Algorithm 1. First, we set the initial set to cyclic permutations of the original sequence (Line 11). If the budget exceeds the current set of output permutations, we iteratively expand the selected permutations by rotating their suffixes (Line 18). This continues till the output set has at least $B$ sequences. Please note that Algorithm 1 is a simplified version of the actual algorithm. For some corner case budget values, we apply certain heuristics to select the right permutations to expand (Line 18). Empirically, we observe that iterative cyclic permutation significantly outperforms random sampling (Figure 8). We conjecture that iterative cyclic rotation generates the most diverse set of $B$ permutations when all $L$ slices have equal deletion probabilities. We will leave the theoretical proof to future research.
Figure 10: An illustration of the sequences selected by the iterative cyclic rotation algorithm for different budgets. We observe that for budget $B\leq L$ the algorithm selects sequences generated by rotating the entire sequence. For budget $L<B\leq L(L-1)$, the algorithm generates newer sequences by rotating the rotating subsequences starting from the second element. This continues as the budget increases and smaller subsequences are rotated.
@@ B.1 Iterative Cyclic Rotation
11:                  $w=\mathrm{score}[o\cup v^{\prime}]$ // compute score for the sequence $o\cup v^{\prime}$ (Eq. 24)
12:                  if $v^{\prime}\notin o$ then $G.\mathrm{add\_edge}(v,v^{\prime},w)$ // check for feasibility
@@
14:         end for
15:         $(V,V^{\prime})=\mathrm{MaximumWeightMatching}(G)$ // use Hungarian algorithm (Kuhn, 1955)
16:         for $o\in\mathcal{O}$ do
17:              // select vertex associated with each sequence from perfect matching
18:              ${v^{*}}=\{v^{\prime}|v^{\prime}\in V^{\prime}\land v=o.\mathrm{pop}()\}$
19:              $o=o\cup{v^{*}}$ // add to sequence
@@
21:     end for
22:     $\mathcal{O}_{B}=\{o\in\mathcal{O}|o\text{ is among top $B$ sequence with highest }\mathrm{score}[o]\}$
23:     return $\mathcal{O}_{B}$
24:end function
B.2 Bipartite-matching based Sequence Selection
Before describing the bipartite selection algorithm, we discuss the scoring function for sequences with given deletion probabilities. Given a slice sequence with deletion probabilities $\mathcal{S}=(p_{1},p_{2},p_{3},p_{4})$, the scoring function computes the expected number of surviving slices after $t$ deletions:
$\mathrm{score}[\mathcal{S},t]=4.(1-p_{1}-p_{2}-p_{3}-p_{4})^{t}+3.(1-p_{1}-p_{2}-p_{3})^{t}+2.(1-p_{1}-p_{2})^{t}+(1-p_{1})^{t}.$ (23)
The above equation is a function of $t$, which needs to be set by the user. In the general case, Eq. 23 can be written as:
$\mathrm{score}[\mathcal{S},t]=\sum_{i=1}^{n}i.\left(1-\sum_{j=1}^{i}p_{j}\right)^{t}.$ (24)
Next, we describe the details of the Bipartite-matching based selection (BMS) algorithm in Algorithm 2. The objective of BMS is to select a set of $B\leq L$ diverse permutation sequences where sequences have a high score (Eq. 24). This algorithm starts by selecting $L$ different starting elements for the sequences in Line 4. Then, BMS iteratively selects the next element within each sequence. This involves constructing a bipartite graph between the last elements of the sequences seen so far and the next set of elements. The edge weights are set to the score of the sequence that is a concatenation of the current sequence, $o$, and the next node, $v^{\prime}$. The edges incident on feasible next elements (elements not seen in a permutation sequence so far) as shown in Line 12. We perform a maximum weight perfect graph matching on the graph, $G$, using the Hungarian algorithm (Kuhn, 1955). Based on the matching, we select the next element for each permutation sequence (Line 18). BMS continues this process till all sequences $o\in\mathcal{O}$ have $L$ elements. Based on the budget, we output $B$ sequences from $\mathcal{O}$ that have the highest sequence scores. For maximum weight perfect matching, (Tomizawa, 1971) provides an efficient algorithm with time complexity $O(n^{3})$, where $n$ is the number of nodes. Using this algorithm, the overall time complexity of BMS is $O(L^{4})$, where $L$ is the number of slices. The following result shows that BMS generates the most diverse sequence set.
Lemma 3.
We also validate the above result empirically and find that BMS produces a range of sequences with high scores (Figure 8). However, it remains an open question whether these outputs are optimal (while considering both diversity and scores).
Figure 11: Illustration of conditional sampling of slice sequences. For each sequence, one slice is sampled at a time based on their deletion probabilities. We observe that conditional sampling selects sequences that has slices with relatively lower deletion probabilities towards the top.
The BMS algorithm can output up to $L$ diverse permutations. Extending it for a budget $B>L$ is non-trivial. Therefore, in this setting, we use a conditional sampling approach to generate diverse sequences. In conditional sampling, for each sequence, we sample slices one at a time based on their deletion probabilities (low deletion probabilities have a higher chance of being sampled). If a sampled sequence already exists in the selected set, it is discarded. An illustration of the conditional sampling approach is shown in Figure 11. We observe that the selected samples (shown in green) have slices with relatively lower deletion probabilities at the top.
B.3 S3T Training Procedure
In this section, we provide an outline for the training procedure within the S3T framework in Algorithm 3. S3T proceeds by dividing the entire dataset into $m$ shards. Each shard is further divided into $L$ disjoint slices. Based on the budget $B$, we obtain the permutation sequences to perform slice-wise training. We train each model $f$ on a unique sequence of slice sequence, $\pi(\mathcal{S})$. We return the complete set of models, $\mathcal{F}$. During inference, the user selects the best-performing model within each shard and deploys an ensemble of those models to production.
@@ B.3 S3T Training Procedure
5:     $\Pi=\mathrm{SelectTrainingSequences}(L,B)$ // using cyclic or BMS algorithm
6:     for $\pi\in\Pi$ do
7:         $\mathcal{S}_{\pi}=\pi(\mathcal{S})$ // order the slices according to the permutation $\pi$
8:         $f=\mathrm{SliceWiseTraining}(\mathcal{S}_{\pi})$
9:         $\mathcal{F}=\mathcal{F}\cup f$
10:     end for
11:end for
12:return $\mathcal{F}$
@@
$\mathcal{F}=\mathcal{F}\cup\overline{\mathcal{F}}$ // introduce updated models
@@ C.1 Experimental Setup
We perform all experiments using PyTorch (Paszke et al., 2019) and Huggingface (Wolf et al., 2019) framework. Our experiments were run on NVIDIA A6000 GPUs. In Table 2, we report the common set of hyperparameters for S3T fine-tuning experiments. All hyperparameters were set using a grid search with the Weights & Biases framework. We use an AdamW optimizer with the corresponding learning rates for each dataset (reported in Table 2). During fine-tuning of the models, we perform full-precision training for all settings except instruction tuning where we use 8-bit training.
@@
C.2 Design Choices
We discuss the rationale behind the design choices for the proposed slice-wise training approach. First, we used top-to-bottom fine-tuning because we found the top layers were easier to train at the start and a bottom-up approach didn’t converge. Note that our training process does not need to proceed in a layer-wise fashion, we can even train multiple LoRA layers per slice. Second, we train layers in a cumulative manner where the $i$-th layer is trained using slices $\{1,\ldots,i\}$. We found that this way of training helps in better convergence compared to the setup where we train every layer using a different slice. Third, we analyze the cost of slice-wise training and find it to be comparable to full training. Considering that training cost is $c=O(nl)$, where $n$ is the dataset size and $l$ is the number of trainable layers (empirical evidence in Appendix C.3). For slice-wise training, we observe that $c=\sum_{i=1}^{k}(\frac{in}{k})(\frac{l}{k})=O\left(\frac{nl}{2}\right)$, which is of the same order as full training.
C.3 Additional Experiments
@@
Performance Degradation with Slice Deletion. In Figure 12, we report the relative performance drop with an increasing number of slices affected by deletion requests. For vision datasets in Figure 12 (left), we only observe a small drop in performance (<5%). For instruction tuning in Figure 12 (middle), we observe a negligible performance drop for most datasets except MMLU & ARC, which are more challenging open-ended QA benchmarks. For text classification in Figure 12 (right), we observe an increased drop in performance for a few of the datasets. We hypothesize that this occurs because the RoBERTaLARGE (used in text classification tasks) is a relatively weaker pre-trained model compared to Llama and ViT. We also observe that the performance can vary based on the task and overall it is dependent on both the task and the model being fine-tuned. For unlearning, a lower performance drop is better as it suggests that we can continue using the same model without retraining for a longer time.
Figure 12: Relative performance drop with an increasing number of deleted slices. We observe a relatively small performance drop for image classification and instruction tuning tasks while noticing a considerable drop for a few of the text classification datasets.
@@
Layer Allocation per Slice. This experiment aims to answer the question: how many slices can we pack into a model while still achieving good performance? We conduct an experiment where we allocate different numbers of PEFT layers per slice and also vary the number of slices. We report the results on the RTE dataset using RoBERTaLARGE in Figure 13. We observe that the performance is poor when the number of layers per slice is low. For example, when the number of layers per slice = 1, the performance improves as we increase the number of slices but again drops when the slice count is too high. This shows that it may not be feasible to train the model using many slices without a drop in performance. Overall, the performance variation depends on the underlying task and model, so the developer should select these hyperparameters based on their performance requirements.
@@
Sequence-based Training. In this experiment, we compare S3T’s performance with sequential training. Sequential training involves training the entire PEFT model on a sequence of slices. This was used in (Kumar et al., 2023) to improve the retraining efficiency of SISA. In Figure 14, we report the performance of S3T and sequential training using ViTBASE on CIFAR-10 and CIFAR-100 datasets. We observe that the performance of S3T gradually increases as it is trained on more slices. Overall, we observe that S3T achieves quite similar or outperforms sequential training. It is important to note that S3T trains a significantly smaller number of parameters ($L$ times reduction compared to sequential training) but still achieves competitive performance.
Figure 15: We compare the scores (Eq. 24) of the generated sequences by BMS and sorted cyclic rotation. We observe BMS consistently outperforms cyclic rotation.
BMS vs. Cyclic rotation with Deletion prior. In this setting, we compare the BMS selection algorithm with a variant of cyclic rotation when the prior deletion probabilities are available. We perform cyclic rotation by first sorting the slices based on their deletion probabilities. For example, we sort the sequence with deletion probabilities: ($S_{1}$: 0.5, $S_{2}$: 0.4, $S_{3}$: 0.1) as: ($S_{3},S_{2},S_{1}$). Then, for a budget $B=3$, the stored sequences are: ($S_{3},S_{2},S_{1}$), ($S_{1},S_{3},S_{2}$), ($S_{2},S_{1},S_{3}$). In this variant, the slices most likely to be deleted are not at the top of any sequences. We follow the experimental setup in Section 4.3 and sample deletion priors using a Dirichlet distribution (over 10 runs). In Figure 15, we report the total sequence scores (Eq. 24) obtained by BMS and sorted cyclic rotation for a budget, $B=L$. We observe that BMS consistently outperforms sorted cyclic rotation for all budgets. Please note that the average edit distance achieved by both methods are the same as cyclic rotations are guaranteed to produce the maximum diversity for budgets: $B\leq L$.
@@
Appendix D Broader Impact
We present a scalable approach to perform exact unlearning in production systems. We hope that this framework will be adopted by organizations and enable a cost-effective way to ensure the privacy requests of various users. In general, unlearning systems are susceptible to attacks where the adversary may design deletion requests in a way to modify the model behavior according to their needs. Therefore, it is important to ensure that the deletion requests executed on the unlearning system are not malicious. However, this is a limitation of unlearning systems in general and not specific to our proposed framework, S3T. Future works can focus on the identification of malicious unlearning requests.